OpenStudy (anonymous):

Solve, using the substitution method. 2x + 4y = 16 x – 7 = – 4y

OpenStudy (arabpride):

2x + 4y = 16 -2x -2x 4y = -2x +16 /4 /4 y = -2/4 + 4 ~~~~~~~~~~~ x - 7 = -4y -x -x -7 = -x -4y +4y +4y 4y - 7 = -x +7 +7 4y = -x + 7 /4 /4 y = x/4 + 7/4 ~~~~~~~~~~~~ HOPE I'm RIGHT!!! I HAVEN'T DONE THIS STUFF IN A WHILE!!

OpenStudy (anonymous):

thanks